OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form

OpenStudy (anonymous):

Equate the slopes and solve for y.\[\frac{y-6}{x-3}=\frac{y+3}{x-0} \]

OpenStudy (anonymous):

y = 3x - 3

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw